OpenStudy (anonymous):

can someone help me, please factor: 36n^2 - 81

OpenStudy (anonymous):

(6n-9)(6n+9) Try it.

OpenStudy (campbell_st):

its the difference of 2 squares a^2 - b^2 = (a - b)(a + b) and 36n^2 = (6n)^2 81 = (9)^2 just plug in the values.

OpenStudy (anonymous):

use this a^2-b^2=(a+b)(a-b) here a^2=36n^2 =>a=6n b^2=81 =>b=9 now put values in above formula and get result.

OpenStudy (anonymous):

thank you all